please help me solve and explian -4(-3m-2)=32.
-17=-2n+13-8n

Answer by checkley79(3341)

-4(-3m-2)=32
12m+8=32
12m=32-8
12m=24
m=24/12
m=2 ans.
-------------------
-17=-2n+13-8n
-17=-10n+13
10n=13+17
10n=30
n=30/10
n=3 ans.